According to Wein’s Law, The peak wavelength of radiation emitted is determined by the temperature of the emitter according to the following equation. λmax=2900/Twhere T is the temperature of the emitter in degrees Kelvin (water freezes at 273 degrees K, 0 degrees Celsius and 32 degrees Fahrenheit). NASA’s Cassini space probe has made several fly-bys of Saturn’s moon Titan, the only moon with an atmosphere. Here is your problem. Given that Earth’s average temperature is 15 degrees Celsius and Titan’s temperature is -180 degrees Celsius, calculate who has the longer peak wavelength (don’t worry about having the correct units) and what the ratio is between the longer and shorter wavelength.
Answer #1

from the Wein's law

λmax=2900/T

on the earth

T = 15+273 = 288 K

λmax=2900 um .K/288 K = 10.06 um

on the Titan’s

T = -180+ 273 =93 K

λmax=2900 um .K/93 K = 31.18 um

Titan has longer wavelength

ratio = 31.18/10.06 =3.09